What is the area of a triangle with the base of 20 inches and the height of 5 inches?
Answer by stanbon(75887)

What is the area of a triangle with the base of 20
inches and the height of 5 inches?
-------------
Area = (1/2)base*height
---
A = (1/2)20*5 = 50 sq inches
===========================
Cheers,
Stan H.
==============